OpenStudy (anonymous):

A wheel of diameter 3.0 cm has a 4.0-m cord wrapped around its periphery. Starting from rest, the wheel is given a constant angular acceleration of 2.0 rad/s^2. The cord will unwind in ?

OpenStudy (anonymous):

circumference of wheel = \(\pi\)d =0.09m No of turns of rope over the circumference = length / circumference ~ 44 The angle subtended by cord = 2\(\pi\) * no of turns = 88\(\pi\) So, the wheel will have to rotate by 88\(\pi\) to unwind the entire rope. There is a kinematical eqn for the rotational motion of the wheel. \[\theta =\omega t+1/2\alpha \ t ^{2}\] where \(\omega\) is the initial angular velocity of wheel \(\alpha\) is the angular acceleration of the wheel. \(\theta\) is the angle displacement of the wheel t is the time of motion Here \(\omega\)=0 \(\alpha\)=2 \(\theta\)=88\(\pi\) Find t

OpenStudy (anonymous):

then t will be ~16 ! Thanx :)

OpenStudy (anonymous):

\[\theta=(lenght/radius) , so \theta=(4/0.015) ~267\] that a simple way to find \[\theta \]